OpenStudy (anonymous):

center and radius of the equation x^2+y^2=81

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2 (h,k) is the center. r is the radius.

hartnn (hartnn):

@chrismae18 Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) The centre and radius of a standard circle \(\Large x^2+y^2 = r^2\) will be Centre = (0,0) and radius = r so what can you say about the radius of your circle ?

OpenStudy (isaiah.feynman):

\[x^{2}+y^{2}=r^{2}\] The radius can be determined by \[\sqrt{x^{2}+y^{2}} =r\]

OpenStudy (isaiah.feynman):